Given P(x) = 2x2 +10x+15 find the function value P(-2) I a) -13 b) 27 c) 3 d) 43 a.. Ob.. O c.. Od.. .
Answer #1

p(x2= - 2x² tontis P(-2)-26-2324100-28415 = -214) +100–2)+15 = -8 – 20115 = -28415 -13 option (A)